A genetic experiment with peas resulted in one sample of offspring that consisted of 421 green peas and 158 yellow peas. 

a. Construct a 90% confidence interval to estimate of the percentage of yellow peas 

b. It was expected that 25% of the offspring peas would be yelow. Given that the percentage of offspring yellow peas is not 25%, do the results contradict expectations? 
Answer #1

x=158

n=421+158=579

p^=x/n

p=158/579=0.273

#90 % CI for proportion is

a) Now to find confidence interval for proportion of yellow peas, the formula is

p) – ze* \/f*(1 – h) <P<ộ+ze* * (1 – ) Vn n

For 90% confidence level , zc = 1.645

(0.273–1.64*1 0.273 + (1 -0.273) 579 <p < 0.273+1.64*1 0.273 * (1 – 0.273) 579

(0.242

#90% confidence interval 0.242

Ansb:

No ,the confidence interval includes so true percentage could easily equal 25%